If 8 men can build a house in 35 days, how many days would it take 14 men.
Answer by TimothyLamb(4379)

r = job/people*time
r = 1/(8*35)
r = 1/(14t)
1/(8*35) = 1/(14t)
14t = 8*35
t = 8*35/14
t = 20 days
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php